OpenStudy (idealist10):

Find the point(s) on the curve of 4x^2+9y^2=36 at which the curvature is largest.

OpenStudy (idealist10):

@amistre64 @zepdrix

OpenStudy (amistre64):

is curvature a representative tangent to a circle at a point?

OpenStudy (idealist10):

\[k(x)=\frac{ \left| y'' \right| }{ [1+(y')^2]^{3/2} }\]

OpenStudy (amistre64):

yeah, then i believe you just fill in the parts

OpenStudy (irishboy123):

by curvature, you mean rate of change of tangent vector with arc length??

OpenStudy (amistre64):

just like a curve has a tangent line, it also has a tangent circle of a given radius

OpenStudy (idealist10):

I need to find the curvature using the formula above, then find the derivative of curvature and set it equal to 0 to find the critical points. And plug the critical points to 4x^2+9y^2=36 to find the y values since the question is asking point(s) where the curvature is largest. But how do I find the curvature first? I'm little stucked on finding the derivatives y' and y".

OpenStudy (amistre64):

osculating (means kissing) circles :)

OpenStudy (amistre64):

well, what is k'?

OpenStudy (idealist10):

You mean I supposed to find the derivative of curvature k'(x) first and then find y' and y" to plug those values into k'(x)?

OpenStudy (amistre64):

well, its simpler to determine k' as is and then input the needed y derivatives

OpenStudy (amistre64):

you have a quotient, and a chain rule

OpenStudy (idealist10):

So what's k'(x)?

OpenStudy (amistre64):

thats the question yes, but moreso k'(y)

OpenStudy (amistre64):

\[k=y''~(1+[y']^2~)^{-3/2}\] \[k'=y'''~(1+[y']^2~)^{-3/2}-\frac{3}{2}~2y'~y''~y''~(1+[y']^2~)^{-5/2}\] \[k'=y'''~(1+[y']^2~)^{-3/2}-3y'~[y'']^2(1+[y']^2~)^{-5/2}\] seems about right to me assuming some tings of course about the absolute values

OpenStudy (idealist10):

Is it really okay to get rid of the absolute values like ^ that?

OpenStudy (amistre64):

if you address it correctly it is

OpenStudy (amistre64):

if y'' -s postive on the interval, then y'' = |y''|

OpenStudy (amistre64):

4x^2+9y^2=36 8x x'+ 18y y'=0 y'= -4/9 x y'' = -4/9 y''' = 0

OpenStudy (amistre64):

|y''| =4/9 so im not worried about it

OpenStudy (idealist10):

8x+18y(dy/dx)=0 18y(dy/dx)=-8x dy/dx=-8x/(18y)=-(4x)/(9y)

OpenStudy (amistre64):

\[k'=y'''~(1+[y']^2~)^{-3/2}-3y'~[y'']^2(1+[y']^2~)^{-5/2}\] \[k'=0~(1+[y']^2~)^{-3/2}+3\frac49x~[\frac49]^2(1+[\frac49x]^2~)^{-5/2}\] \[0=3\frac49x~[\frac49]^2(1+[\frac49x]^2~)^{-5/2}\] when x=0

OpenStudy (idealist10):

But how did you get y'=-4/9 x when I got dy/dx=-(4x)/9y?

OpenStudy (amistre64):

-4/9 x = -4x/(9y)

OpenStudy (amistre64):

i see, i dropped a y didnt i

OpenStudy (amistre64):

y is still in the denominator ... so what zeros out is the top part

OpenStudy (amistre64):

4x^2+9y^2=36 an ellipse centered at the origin (0,+-2) (+-3,0) 

OpenStudy (amistre64):

we would expect the largest center of the curve around an ellipse to have the biggest curvature right?

OpenStudy (amistre64):

\[0=3\frac49\frac{x}{\color{red}y}~[\frac49]^2(1+[\frac{4x}{9y}]^2~)^{-5/2}\] \[0=\frac{3.4.4^2~x}{9.9^2.y.(1+[\frac{4x}{9y}]^2~)^{-5/2}}\] it still equals zero when x=0

OpenStudy (anonymous):

@amistre64 it looks like you're saying the curvature is greatest at the points on the y-axis. (Please correct me if I'm wrong, I only read the last few posts.) Here's a quick animation made in Mathematica that demonstrates the curvature is highest at the points along the major axis (on the x-axis). http://imgur.com/78JtIpA

OpenStudy (amistre64):

yeah that is what was going thru my head at the time. i might have been confusing the largest osculating circle with the 'fastest' rate of curve tho in hindsight. thnx for the insight :)